# Outcomes truncated by death in RCTs: a simulation study on the survivor average causal effect

### v2

@@ 1 Introduction
An example of such an RCT is the EpoRepair trial, an ongoing placebo-controlled, parallel group, double-blind RCT on the effect of erythropoietin (Epo) for the repair of cerebral injury in very preterm infants suffering from intraventricular hemorrhage (Rüegger et al., , 2015; Wellmann et al., , 2022). The primary outcome is the composite intelligence quotient (IQ) at five years of age measured using the Kaufmann Assessment Battery for Children (K-ABC). Cognitive development at two years of age, measured using the subscore for cognition of the Bayley Scales of Infant and Toddler Development (BSID-III), is a secondary outcome. Both, IQ and cognitive development are measured only once. The EpoRepair trial randomized 121 preterm infants to Epo or Placebo in a 1:1 ratio. However, 19 (15.7 %) of these vulnerable infants died before the age of two years (15 died before term equivalent age). How should the effect of the intervention (Epo) on IQ or on cognitive development be estimated in this case, or more conceptually, which estimand should be targeted, when a considerable proportion of outcome measurements is truncated by death?
@@
One commonly used ad hoc solution to the problem is to restrict the analysis to the survivors (if all survivors have a measurement of the outcome) or more generally, to complete cases. In fact, the use of complete case analysis in RCTs with missing data is prevalent (Zhang et al., , 2017; Ren et al., , 2022), although it is well-known that this leads to biased estimates of the treatment effect, as it corresponds to a nonrandomized comparison (see for example CHMP, , 2010; Little et al., , 2012). No bias would occur only when outcome measurements are “missing completely at random”, i.e., the missingness is neither related to the assigned treatment nor to any characteristics of the patients. The size of the bias depends on the amount of missing or truncated observations and on their relation with the assigned treatment. Such a relationship with the patients response to the treatment is often likely, and differing proportions of missing observations between trial arms may add to this concern (Vickers and Altman, , 2013). In situations when outcome measurements are not “missing completely at random”, the treatment effect estimated by complete case analysis usually does not relate to a clear or relevant clinical question.
@@
Another ad hoc solution may be to treat the truncated outcomes as missing outcome measurements and use multiple imputation to recover the missing information. The underlying assumption of multiple imputation is that measurements are “missing at random”, which means that the missingness can be explained by observed data. Multiple imputation has been recommended for analyzing trials with missing data (but recommendations are not restricted to trials) and is known to be a better choice than single imputation methods (White et al., , 2011; Groenwold et al., , 2012; Vickers and Altman, , 2013). A strength of multiple imputation is that the uncertainty about the imputations can be taken into account using Rubin’s rules (Rubin, , 1987) to combine the estimates from the imputed data sets. With the availability of software to generate imputations and to combine effect size estimates, for instance in the R packages mice (van Buuren and Groothuis-Oudshoorn, , 2011) or Amelia (Honaker et al., , 2011) this method has become increasingly popular. Multiple imputation is certainly a good choice for imputing missing outcomes if they would have been observable. However, outcome measurements truncated by death are not just a missing data problem: they are not defined (the IQ of a child who died before the IQ could be measured is not defined). While using multiple imputation allows to analyze all randomized patients (including those who died) the estimated causal treatment effect is “hypothetical”, and for a scenario in which no patient would have died. Such a “hypothetical strategy” is usually more meaningful if the intercurrent event (unlike death) could be avoided by design in a future trial. Nevertheless, in practice multiple imputation may sometimes be used to analyze trials with outcomes truncated by death, especially if truncation by death accounts for only part of the missing outcomes. Multiple imputation may then also be used to impute missing outcomes of survivors only, which may be of limited use if some observations are truncated by death. For the EpoRepair trial, it was planned that multiple imputation will be used to adjust for missing outcomes due to death or dropout (Rüegger et al., , 2015). We therefore focus on multiple imputation to address a hypothetical estimand, although there are alternatives like inverse probability weighting or G-computation (see Olarte Parra et al., , 2023, who are linking these methods).
@@
Despite these reservations, we believe that the SACE is a useful estimand in the setting of the EpoRepair trial, for several reasons. First, the principal stratum strategy should be more relevant than the hypothetical strategy if an intercurrent event in a specific population cannot be avoided by design, as it is unfortunately the case with death among very preterm infants (and in other vulnerable or severely ill populations). Second, the majority of infants survived past 2 years, and since a treatment effect on survival was not a priori expected, we expect the majority of the survivors to be always survivors (so we expect the principal stratum exists and even includes the majority of infants).
@@ 2.3 Estimands
The estimates derived by the three methods (see Section 2.4) were compared in terms of bias, mean squared error and coverage with regard to two estimands. The first estimand $\theta_{1}$ is the treatment effect on the outcome used in the simulation, equivalent to $b_{z}$, corresponding to the average causal effect $E[Y(1)-Y(0)]$. $\theta_{1}$ is a hypothetical estimand in the presence of mortality and represents the causal effect of the treatment on the outcome in the absence of mortality. The second estimand $\theta_{2}$ is the survivor average causal effect (SACE), the treatment effect on the “always survivors”, the patients who would have survived under both treatments, defined as $E[Y(1)-Y(0)\,|\,S(1)=S(0)=1]$. $\theta_{2}$ is a principal stratum estimand and is relevant here, because mortality cannot be eliminated in the population at hand, and because a treatment effect on survival cannot be ruled out. While $\theta_{1}$ is clearly a marginal estimand, $\theta_{2}$ may be seen as a marginal estimand for the subpopulation of “always survivors” or as a conditional expectation regarding survival status (McGuinness et al., , 2019; Luo et al., , 2022). Note that the average treatment effect on survivors, which is estimated by complete case analysis, is defined as $E[Y(1)\,|\,S(1)=1]-E[Y(0)\,|\,S(0)=1]$.
@@
Here, $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes of patient $i$ on treatment and control, and $S_{i}(1)$ and $S_{i}(0)$ are the potential survival outcomes of patient $i$ under treatment and control, respectively (if $S_{i}(1)=1$, patient $i$ survived on treatment, if $S_{i}(1)=0$, patient $i$ died on treatment; if $S_{i}(0)=1$ patient $i$ survived on control, if $S_{i}(0)=0$ patient $i$ died on control). To have one estimand per scenario, we then averaged over all simulations to derive $\theta_{2}$ as $\frac{1}{n_{\text{sim}}}\sum\limits_{s=1}^{n_{\text{sim}}}\theta_{2s}$. Note that $\theta_{2}$ is strictly speaking an estimate of the SACE but serves as the estimand in our simulation study. An alternative method to derive $\theta_{2}$ that was listed in the simulation study protocol was omitted, since it was less intuitive and results were very similar.
@@ 2.4 Methods to be evaluated
where $i$ indexes over patients assigned to arm $z$ = 1 and $j$ indexes over patients assigned to arm $z$ = 0. The estimates $\hat{p}_{i}(0)$ and $\hat{p}_{j}(1)$ refer to the probability of survival under Placebo and Epo, as estimated by logistic regression from separate fits to the Placebo and Epo arms, respectively. Complete case analysis was performed on the survivors only and is used as a benchmark for a biased analysis in most situations, since the treatment effect on survivors is not a randomized comparison. Complete case analysis was done using a linear model with treatment as explanatory variable. Only in the absence of mortality or if mortality occurred “completely at random” would complete case analysis provide an unbiased estimator for both, $\theta_{1}$ and $\theta_{2}$, which then coincide with the treatment effect on survivors. The treatment effect estimate derived by analysis of multiply imputed data can be expected to be close to $\theta_{1}$, independent of the treatment effect on survival. Complete case analysis and the SACE estimator do not conform with the intention-to-treat (ITT) principle, as only a subset of patients is analyzed. In contrast, multiple imputation conforms with the ITT principle but creates unobservable (hypothetical) data.
@@
As previously mentioned, the SACE is not identifiable without strong assumptions. A comprehensive overview of principal stratification methods and their underlying assumptions is given in Lipkovich et al., (2022). Hayden et al., make the stable unit treatment value assumption (SUTVA, Rubin, , 1980) and the explainable nonrandom survival assumption (Robins, , 1998; Hayden et al., , 2005, adapted from explainable nonrandom noncompliance). The SUTVA implies that a subject’s observed outcome is the same as (consistent with) the potential outcome associated with the treatment that the subject was assigned/randomized to and that the subject’s potential outcomes do not depend on (or interfer with) the treatment assigned to other subjects (Bornkamp et al., , 2021; Lipkovich et al., , 2022). The SUTVA needs to be made by most principal stratification methods (and other causal inference approaches), since it allows to connect potential and observed outcomes at an individual patient level. One reason for choosing Hayden’s SACE estimator was that it neither relies on the monotonicity assumption nor on the exclusion restriction assumption (two other common assumptions), which are both unrealistic in the context of the SACE and the EpoRepair study. Monotonicity would imply that patients who die on the experimental treatment would also die on the control treatment, i.e., that there are no control-only survivors. Exclusion restriction may be reasonable when the aim is to estimate the CACE, the average treatment effect on compliers, defined as those who would take the intervention when randomized to intervention and control when randomized to control both throughout the trial (e.g., Table 1 in Lipkovich et al., , 2022). Assuming exclusion restriction in this setting would imply that potential outcomes for never-takers (those who never take the intervention) and always-takers (those who always take the intervention) are the same, regardless of the randomized treatment, as they actually take the same treatment. In contrast, assuming exclusion restriction for the SACE would imply that potential outcomes of always-survivors and never-survivors (sometimes called the doomed) would be the same. This would be an unrealistic assumption, as it would imply no treatment effect in this stratum for which we want to estimate one. Instead, Hayden’s SACE estimator relies on modeling to identify strata membership based on observed baseline covariates, assuming explainable nonrandom survival:
@@
The first assumption states that, conditional on the baseline covariates $X_{i}$, the survival status $S_{i}(z)$ of a subject under treatment $z$ is independent of its survival status $S_{i}(1-z)$ under treatment $1-z$. The second assumption states that, conditional on surviving when assigned to treatment $1-z$, and on the baseline covariates $X_{i}$, the survival status $S_{i}(z)$ of a subject under treatment $z$ is independent of its outcome $Y_{i}(1-z)$ under treatment $1-z$. These assumptions essentially mean that no unmeasured confounders are present. However, since survival under treatment $z$ and $1-z$ can never be jointly observed, explainable nonrandom survival may be called a “cross-world” assumption, which is not testable on the data. In summary, explainable nonrandom survival is a strong and untestable assumption (Lipkovich et al., , 2022; Kurland et al., , 2009), but in the case of the SACE may still be less problematic than other assumptions. Further, in order to make the assumption more plausible, one should collect and use baseline covariates which are strongly predictive for survival (which we tried), and ideally incorporate an analysis of sensitivity of results to departures from Equations (7) and (8).
@@
Multiple imputation of missing outcomes was performed using the R package mice (van Buuren and Groothuis-Oudshoorn, , 2011), generating 10 imputations per missing value. Similar to the analysis using the SACE estimator, we once used all covariates that were used in the simulation of patient outcomes (gestational age, head circumference and SES) as predictors for the imputation of the missing outcome values, and omitted the covariate head circumference or gestational age in two additional analyses. Here, the omission of these covariates should lead to a violation of the “missing at random” assumption. The analysis model was the same linear model used for complete case analysis, just applied to the imputed data. Results were pooled across all imputations using Rubin’s rules, as implemented in mice.
@@ 3.2 Treatment effect estimates and bias
Figure 1 shows the average estimates of the treatment effect on the outcome for each scenario and method together with the estimands $\theta_{1}$ and $\theta_{2}$. For the SACE estimator and for multiple imputation, the average estimates from the analysis using all covariates (green and red) are shown together with those from the analyses omitting either of the covariates head circumference or gestational age (two lighter shades of green and red). Across all scenarios the two estimands, the causal effect of the treatment in the absence of mortality ($\theta_{1}$) and the survivor average causal effect ($\theta_{2}$), were similar. In scenarios without a treatment effect on survival (B, E, H, middle row of Figure 1), treatment effect estimates were similar for all methods and close to the causal effect used in the simulation. In particular, in the absence of a treatment effect on both outcome and survival (E) all methods estimated very similar treatment effects, even in presence of a significant proportion of outcomes truncated by death.
@@
As already indicated in Figure 1, the average bias of the SACE estimator and multiple imputation with regard to both $\theta_{1}$ and $\theta_{2}$ was much lower, compared to the bias of complete case analysis, when all covariates were used or when head circumference was omitted. Omitting gestational age considerably increased the bias of the SACE estimator and multiple imputation, with 95 % confidence intervals that exclude a bias of 00, but the bias of these analyses was still much smaller than that of complete case analysis. Since the SACE estimator targets $\theta_{2}$ (estimated as $\hat{\theta}_{2}$), and multiple imputation targets $\theta_{1}$, we expected that the corresponding bias would be lower for these combinations of method and estimand than for the others. However, although the average bias of the SACE estimator with regard to $\theta_{2}$ and of multiple imputation with regard to $\theta_{1}$ was small in all scenarios (Table 4), and 95 % confidence intervals for bias included 0, except for analyses omitting gestational age (Figure 2), the expected patterns were not consistently observed. The summary Table 4 indicates that multiple imputation was not generally less biased with regard to $\theta_{1}$ than $\theta_{2}$. Particularly in scenarios with a negative treatment effect on survival (rows with OR=0.5 in Table 4), this pattern was sometimes even reversed. This may be due to the larger mortality in these scenarios which resulted in smaller numbers of patients analyzed by the SACE estimator and larger number of missing values which were multiply imputed. With regard to $\theta_{2}$, multiple imputation was on often more biased than the SACE estimator. Further, the analyses omitting the covariate head circumference or gestational age with the SACE estimator resulted in slightly larger and much larger average bias with regard to $\theta_{2}$ than the corresponding analysis with all covariates, and the same applied for multiple imputation with regard to $\theta_{1}$.

### v3

@@ 1 Introduction
An example of such an RCT is the EpoRepair trial, an ongoing placebo-controlled, parallel group, double-blind RCT on the effect of erythropoietin (Epo) for the repair of cerebral injury in very preterm infants suffering from intraventricular hemorrhage (Rüegger et al., , 2015; Wellmann et al., , 2022). The primary outcome is the composite intelligence quotient (IQ) at five years of age measured using the Kaufmann Assessment Battery for Children (K-ABC). Cognitive development at two years of age, measured using the subscore for cognition of the Bayley Scales of Infant and Toddler Development (BSID-III), is a secondary outcome. Both, IQ and cognitive development are measured only once. The EpoRepair trial randomized 121 preterm infants to Epo or Placebo in a 1:1 ratio. However, 19 (15.7 %) of these vulnerable infants died before the age of two years (15 died before term equivalent age). Further, a group blinded interim safety analysis indicated an imbalance in mortality between trial arms. How should the effect of the intervention (Epo) on IQ or on cognitive development be estimated in this case, or more conceptually, which estimand should be targeted, when a considerable proportion of outcome measurements is truncated by death?
@@
One commonly used ad hoc solution to the problem is to restrict the analysis to the survivors (if all survivors have a measurement of the outcome) or more generally, to complete cases. In fact, the use of complete case analysis in RCTs with missing data is prevalent (Zhang et al., , 2017; Ren et al., , 2022), although it is well-known that this leads to biased estimates of the treatment effect, e.g. with regard to a treatment policy estimand, as it corresponds to a nonrandomized comparison (see for example CHMP, , 2010; Little et al., , 2012). No bias would occur only when outcome measurements are “missing completely at random”, i.e., the missingness is neither related to the assigned treatment nor to any characteristics of the patients. The size of the bias depends on the amount of missing or truncated observations and on their relation with the assigned treatment. Such a relationship with the patients response to the treatment is often likely, and differing proportions of missing observations between trial arms may add to this concern (Vickers and Altman, , 2013). In situations when outcome measurements are not “missing completely at random”, the treatment effect estimated by complete case analysis usually does not relate to a clear or relevant clinical question. This implies that it will also be biased with regard to estimands other than the treatment policy estimand.
@@
Another ad hoc solution may be to treat the truncated outcomes as missing outcome measurements and use multiple imputation to recover the missing information. The underlying assumption of multiple imputation is that measurements are missing at random, which means that the missingness can be explained by observed data. Multiple imputation has been recommended for analyzing trials with missing data (but recommendations are not restricted to trials) and is known to be a better choice than single imputation methods (White et al., , 2011; Groenwold et al., , 2012; Vickers and Altman, , 2013). A strength of multiple imputation is that the uncertainty about the imputations can be taken into account using Rubin’s rules (Rubin, , 1987) to combine the estimates from the imputed data sets. With the availability of software to generate imputations and to combine effect size estimates, for instance in the R packages mice (van Buuren and Groothuis-Oudshoorn, , 2011) or Amelia (Honaker et al., , 2011) this method has become increasingly popular. Multiple imputation is certainly a good choice for imputing missing outcomes if they would have been observable. However, outcome measurements truncated by death are not just a missing data problem: they are not defined (the IQ of a child who died before the IQ could be measured is not defined). While using multiple imputation allows to analyze all randomized patients (including those who died) the estimated causal treatment effect is “hypothetical”, and for a scenario in which no patient would have died. Such a “hypothetical strategy” is usually more meaningful if the intercurrent event (unlike death) could be avoided by design in a future trial. Nevertheless, in practice multiple imputation may sometimes be used to analyze trials with outcomes truncated by death, especially if truncation by death accounts for only part of the missing outcomes. Multiple imputation may then also be used to impute missing outcomes of survivors only, which may be of limited use if some observations are truncated by death. For the EpoRepair trial, it was planned that multiple imputation will be used to adjust for missing outcomes due to death or dropout (Rüegger et al., , 2015). We therefore focus on multiple imputation to address a hypothetical estimand, although there are alternatives like inverse probability weighting or G-computation (see Olarte Parra et al., , 2023, who are linking these methods).
@@
Despite these reservations, we believe that the SACE is a useful estimand in the setting of the EpoRepair trial, for several reasons. First, the principal stratum strategy should be more relevant than the hypothetical strategy if an intercurrent event in a specific population cannot be avoided by design, as it is unfortunately the case with death among very preterm infants (and in other vulnerable or severely ill populations). Second, while the majority of infants was expected to survive past 2 years, no treatment effect on survival was expected a priori. As a consequence, we can expect the majority of the survivors to be always survivors (so we expect the principal stratum exists and even includes the majority of infants).
@@ 2.3 Estimands
The estimates derived by the three methods (see Section 2.4) were compared in terms of bias, mean squared error and coverage with regard to two estimands. The first estimand $\theta_{1}$ is the treatment effect on the outcome used in the simulation, equivalent to $b_{z}$ (see Section 2.2), corresponding to the average causal effect $E[Y(1,\text{survival until 2 years})-Y(0,\text{survival until 2 years})]$. $\theta_{1}$ is a hypothetical estimand and represents the causal effect of the treatment on the outcome in the absence of mortality (up to 2 years of age). The second estimand $\theta_{2}$ is the survivor average causal effect (SACE), the treatment effect on the “always survivors”, the patients who would have survived under both treatments, defined as $E[Y(1)-Y(0)\,|\,S(1)=S(0)=1]$. $\theta_{2}$ is a principal stratum estimand and is relevant here, because mortality cannot be eliminated in the population at hand, and because a treatment effect on survival cannot be ruled out. While $\theta_{1}$ is clearly a marginal estimand, $\theta_{2}$ may be seen as a marginal estimand for the subpopulation of “always survivors” or as a conditional expectation regarding survival status (McGuinness et al., , 2019; Luo et al., , 2022). Note that the average treatment effect on survivors, which is estimated by complete case analysis, is defined as $E[Y(1)\,|\,S(1)=1]-E[Y(0)\,|\,S(0)=1]$.
@@
Here, $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes of patient $i$ on treatment and control, and $S_{i}(1)$ and $S_{i}(0)$ are the potential survival outcomes of patient $i$ under treatment and control, respectively (if $S_{i}(1)=1$, patient $i$ survived on treatment, if $S_{i}(1)=0$, patient $i$ died on treatment; if $S_{i}(0)=1$ patient $i$ survived on control, if $S_{i}(0)=0$ patient $i$ died on control). To have one estimand per scenario, we then averaged over all simulations to derive $\theta_{2}$ as $\frac{1}{n_{\text{sim}}}\sum\limits_{s=1}^{n_{\text{sim}}}\theta_{2s}$. The resulting values of $\theta_{2}$ are shown in Figure 1. Note that $\theta_{2}$ is strictly speaking an estimate of the SACE but serves as the estimand in our simulation study. An alternative method to derive $\theta_{2}$ that was listed in the simulation study protocol was omitted, since it was less intuitive and results were very similar.
@@ 2.4 Methods to be evaluated
where $i$ indexes over patients assigned to arm $z$ = 1 and $j$ indexes over patients assigned to arm $z$ = 0. The estimates $\hat{p}_{i}(0)$ and $\hat{p}_{j}(1)$ refer to the probability of survival under Placebo and Epo, as estimated by logistic regression from separate fits to the Placebo and Epo arms, respectively. Complete case analysis was performed on the survivors only and is used as a benchmark for a biased analysis in most situations, since the treatment effect on survivors is not a randomized comparison. Complete case analysis was done using a linear model with treatment as explanatory variable. Only in the absence of mortality or if mortality occurred “completely at random” would complete case analysis provide an unbiased estimator for both, $\theta_{1}$ and $\theta_{2}$, which then coincide with the treatment effect on survivors. Given the underlying missing at random assumption is reasonable and that multiple imputation can make use of available characteristics that are sufficiently associated with the outcome to be imputed, the treatment effect estimate derived by analysis of multiply imputed data can be expected to be close to $\theta_{1}$, independent of the treatment effect on survival. Complete case analysis and the SACE estimator do not conform with the intention-to-treat (ITT) principle, as only a subset of patients is analyzed. In contrast, multiple imputation conforms with the ITT principle but creates unobservable (hypothetical) data.
@@
As previously mentioned, the SACE is not identifiable without strong assumptions. A comprehensive overview of principal stratification methods and their underlying assumptions is given in Lipkovich et al., (2022). Hayden et al., make the stable unit treatment value assumption (SUTVA, Rubin, , 1980) and the explainable nonrandom survival assumption, which they defined themselves and named it in the style of explainable nonrandom noncompliance (Robins, , 1998), despite some important differences (see Vansteelandt and Van Lancker, , 2024). The SUTVA implies that a subject’s observed outcome is the same as (consistent with) the potential outcome associated with the treatment that the subject was assigned/randomized to and that the subject’s potential outcomes do not depend on (or interfer with) the treatment assigned to other subjects (Bornkamp et al., , 2021; Lipkovich et al., , 2022). The SUTVA needs to be made by most principal stratification methods (and other causal inference approaches), since it allows to connect potential and observed outcomes at an individual patient level. One reason for choosing Hayden’s SACE estimator was that it neither relies on the monotonicity assumption nor on the exclusion restriction assumption (two other common assumptions), which are both unrealistic in the context of the SACE and the EpoRepair study. Monotonicity would imply that patients who die on the experimental treatment would also die on the control treatment, i.e., that there are no control-only survivors. Exclusion restriction may be reasonable when the aim is to estimate the CACE, the average treatment effect on compliers, defined as those who would take the intervention when randomized to intervention and control when randomized to control both throughout the trial (e.g., Table 1 in Lipkovich et al., , 2022). Assuming exclusion restriction in this setting would imply that potential outcomes for never-takers (those who never take the intervention) and always-takers (those who always take the intervention) are the same, regardless of the randomized treatment, as they actually take the same treatment. In contrast, assuming exclusion restriction for the SACE would imply that potential outcomes of always-survivors and never-survivors (sometimes called the doomed) would be the same. This would be an unrealistic assumption, as it would imply no treatment effect in this stratum for which we want to estimate one. Instead, Hayden’s SACE estimator relies on modeling to identify strata membership based on observed baseline covariates, assuming explainable nonrandom survival:
@@
The first assumption states that, conditional on the baseline covariates $X_{i}$, the survival status $S_{i}(z)$ of a subject under treatment $z$ is independent of its survival status $S_{i}(1-z)$ under treatment $1-z$. The second assumption states that, conditional on surviving when assigned to treatment $1-z$, and on the baseline covariates $X_{i}$, the survival status $S_{i}(z)$ of a subject under treatment $z$ is independent of its outcome $Y_{i}(1-z)$ under treatment $1-z$. However, since survival under treatment $z$ and $1-z$ can never be jointly observed, explainable nonrandom survival may be called a “cross-world” assumption, which is not testable on the data. In summary, explainable nonrandom survival is a strong and untestable assumption (Lipkovich et al., , 2022; Kurland et al., , 2009). Hayden et al., (2005) suggested and performed a sensitivity analysis for departures from assumptions (7) and (8), which they conducted on data from the ADRSnet clinical trial. They concluded that the significant treatment effect on days to return home (secondary outcome in ADRSnet, only observed in survivors) is fairly robust to departures from the independence assumptions, except in the presence of strong unappreciated interactions between covariates and treatment. While Qu et al., (2023) also concluded that assumptions (7) and (8) may be less implausible than monotonicity, these assumption are among those scrutinized by Vansteelandt and Van Lancker, (2024), who provide a thorough discussion of assumptions made in order to address principal stratum estimands, with a focus on treatment adherence (i.e., the CACE, since adherence and compliance are often used interchangeably). However, in the case of the SACE, if we cannot assume exclusion restriction and do not want to assume monotonicity, we cannot avoid making the above or similar assumptions. Further, in order to make the explainable nonrandom survival assumption more plausible, one should collect and use baseline covariates which are strongly predictive for survival (which we tried), and ideally incorporate an analysis of sensitivity of results to departures from Equations (7) and (8).
@@
Multiple imputation of missing outcomes was performed using the R package mice (van Buuren and Groothuis-Oudshoorn, , 2011), generating 10 imputations per missing value. Similar to the analysis using the SACE estimator, we once used all covariates that were used in the simulation of patient outcomes (gestational age, head circumference and SES) as predictors for the imputation of the missing outcome values, and omitted the covariate head circumference or gestational age in two additional analyses. Here, the omission of these covariates should lead to a violation of the missing at random assumption. The analysis model was the same linear model used for complete case analysis, just applied to the imputed data. Results were pooled across all imputations using Rubin’s rules, as implemented in mice.
@@ 3.2 Treatment effect estimates and bias
Figure 1 shows the average estimates of the treatment effect on the outcome for each scenario and method together with the estimands $\theta_{1}$ and $\theta_{2}$. For the SACE estimator and for multiple imputation, the average estimates from the analysis using all covariates (green and red) are shown together with those from the analyses omitting either of the covariates head circumference or gestational age (two lighter shades of green and red). Across all scenarios the two estimands, the causal effect of the treatment in the absence of mortality ($\theta_{1}$) and the survivor average causal effect ($\theta_{2}$), were similar. In scenarios without a treatment effect on survival (B, E, H, middle row of Figure 1), treatment effect estimates were similar for all methods and close to the causal effect used in the simulation. In particular, in the absence of a treatment effect on both outcome and survival (E) all methods estimated very similar treatment effects, even in presence of a significant proportion of outcomes truncated by death. Since all estimands are equivalent under these conditions, this result could be expected.
@@
As already indicated in Figure 1, the average bias of the SACE estimator and multiple imputation with regard to both $\theta_{1}$ and $\theta_{2}$ was much lower, compared to the bias of complete case analysis, when all covariates were used or when head circumference was omitted. Omitting gestational age considerably increased the bias of the SACE estimator and multiple imputation, with 95 % confidence intervals that exclude a bias of 00, but the bias of these analyses was still much smaller than that of complete case analysis. Since the SACE estimator targets $\theta_{2}$ (estimated as $\hat{\theta}_{2}$), and multiple imputation targets $\theta_{1}$, we expected that the corresponding bias would be lower for these combinations of method and estimand than for the others. However, although the average bias of the SACE estimator with regard to $\theta_{2}$ and of multiple imputation with regard to $\theta_{1}$ was small in all scenarios (Table 4), and 95 % confidence intervals for bias included 0, except for analyses omitting gestational age (Figure 2), the expected patterns were not consistently observed. The summary Table 4 indicates that multiple imputation was not generally less biased with regard to $\theta_{1}$ than $\theta_{2}$. Particularly in scenarios with a negative treatment effect on survival (rows with OR=0.5 in Table 4), this pattern was sometimes even reversed. This may be due to the larger mortality in these scenarios which resulted in smaller numbers of patients analyzed by the SACE estimator and larger number of missing values which were multiply imputed. With regard to $\theta_{2}$, multiple imputation was often more biased than the SACE estimator. Further, the analyses omitting the covariate head circumference or gestational age with the SACE estimator resulted in slightly larger and much larger average bias with regard to $\theta_{2}$ than the corresponding analysis with all covariates, and the same applied for multiple imputation with regard to $\theta_{1}$. The observed increase in bias could be expected due to violation of the explainable nonrandom survival assumption relevant for $\theta_{2}$ and the missing at random assumption relevant for $\theta_{1}$.